OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) 3x – 6y = 12 9x + 2y = –24

OpenStudy (anonymous):

i believe you can graph it by hand after finding points for each equation Now just solving for y to make equation to the form y=mx + b linear 3x -6y=12 => y= (3x -12)/6= x/2 -2 9x + 2y =24 => y= =(-9x +24)/2= -9x/2 +12 Graph each equation by finding 2 coordinate points for x, just make up 2 x values and finding their y's and graph in each. After you graph, 2 lines will intersect at one point...and that point should be the solution of your system.... Just some hints at first, hope this helps